# Harvesting Private Medical Images in Federated Learning Systems with Crafted Models

### v1

@@ V-B Reconstruction Results
In Tab. I we demonstrate the performance of MediLeak over different recovery batch sizes (i.e. the number of samples recovered simultaneously held by the target victim) ranging from 100 to 500 images. We can observe that both the recovery rate (i.e. the ratio of successfully recovered images) and the quantitative scores (i.e. the PSNR and SSIM scores) decrease when the recovered batch size increases. This is expected because the larger the recovery batch size, the more difficult the recovery task to conduct. But in general, MediLeak achieves high recovery rates ($>0.8$) and quantitative scores (PSNR$>80$ and SSIM$>0.9$) for both datasets under all recovery batch sizes. Particularly, the SSIM scores (ranging from 0 to 1) remain to be 0.99 for all settings, because of the recovery excellency. This can be further verified by the recovered samples we visualized in Fig. 3, in which we plot the original images in the first row and the recovered ones in the second. We find that the recovered images are of high quality and cannot be visually distinguished from the original ones, even for some detailed small marks and notations. In Tab. I we also demonstrate the attack time (in seconds). We find that the attack time is monotonically increasing with respect to the recovery batch size. For the largest batch size (i.e. 500 images) over the complex COVIDx CXR-4 dataset, it only takes the attacker less than 10 seconds to fulfill the recovery task, indicating that MediLeak is very effective.
V-C Benchmark Comparison
@@
We compare MediLeak’s attack performance with three optimization-based model inversion attacks (MIAs) including the DLG/iDLG [20, 21], InvertGradient [22], and GradInversion [23] attacks (denoted as DLG, Invert, and GradInvert respectively) over the MedMNIST dataset for one small batch of input images. We compare the PSNR scores, SSIM scores, and the attack time between the three attacks and our work. The results are demonstrated in Fig. 4. We can find that our attack achieves much better PSNR scores and SSIM scores than the existing MIAs, indicating that our attack can reconstruct samples with better quality. At the same time, our attack consumes significantly less time than the current MIAs. Particularly, the existing attacks consume a few hundred seconds to reconstruct one batch of samples, while our attack only requires less than one second, which reduces the current cost by two orders of magnitude. The reason why our attack is much more efficient is that our attack only involves closed-form mathematical calculations while the other three attacks require costly iterative-based optimization methods. However, there is no free lunch and we clarify that the three benchmark works adopt an honest-but-curious attack model, which does not allow the attacker to modify the model parameters and architecture as we did.
V-D Performance Affecting Factors
As we have discussed at the end of section IV. The neuron number $k$ of the first linear layer is the key factor that affects the performance of MediLeak. In this section, we fix the reconstruction batch size to 100 to evaluate the attack performance under different $k$ settings. We demonstrate the recovery rates and the quantitative scores (i.e. the PSNR score and SSIM score) of the successfully recovered samples. For the recovery rate, we observe that it is monotonically increasing with $k$ and obtains relatively high recovery rates when $k\geq m$ (m=100 in this experiment). This is consistent with our theoretical analysis as more bins (larger $k$) will decrease the recovery conflict probability and increase the recovery success rates. For the PSNR and SSIM scores, we observe that the successful samples continuously maintain excellent reconstruction quality (i.e. obtaining very high scores) under different settings, while the unsuccessful ones achieve poor scores lower than the threshold (i.e. $PSNR<20$ and $SSIM<0.9$). This indicates that our attack demonstrates a binary recovery property, meaning that the recovery task either fails to obtain nothing or succeeds in recovering images with perfect quality.
V-E Downstream Tasks
@@
To further evaluate the performance of our attack on clinically relevant downstream tasks, we perform a binary disease classification (the detection of COVID-19) task on both the recovered samples and the actual samples. We use the state-of-the-art vision transformer model (ViT-S) (embedding size=368, number of heads=6, 22M parameters) pre-trained by self-supervised learning (SSL) technique on 30k COVIDx CXR-3 samples and fine-tuned on the RSNA-RICORD part of the dataset to perform the classification task [39] and evaluate it on the COVIDx CXR-4 dataset. We demonstrate the performance in Tab.II. We use widely used machine learning metrics to evaluate the classification performance and we find that the recovered images achieve nearly the same performance as the original ones. This shows that our reconstruction process is highly successful in keeping all semantic meaning within the images and the reconstructed images can be used to perform any potential clinical analysis, which further indicates the severity of the privacy threat imposed by our attack. We consider it a very practical attack scenario for a curious party, which either can be the medical federated learning’s participants or a third-party service provider (who provides the necessary platform, computation resource, and other FL infrastructures) to launch our attack to first reconstruct the sensitive medical images and then feed them to certain downstream analysis tasks to obtain further information about of the victims.
VI Discussion and Future Work
@@
Auxiliary Dataset: The number and quality of samples in the auxiliary data $D_{aux}$ affect the performance of our attack. The dataset is used in the attack preparation phase for estimating essential attack parameters before crafting the “zero gradient” and “’linear leakage’ modules. In the ideal case, the auxiliary dataset shall have the same data distribution as the target victim’s local dataset or the auxiliary dataset is more representative. It may be a challenge for the general vision tasks to find such a representative auxiliary dataset. However, for the medical images, this does not pose a barrier because the medical images such as the CT scanning images of humans are fixed to have the same data format and similar common features. The attacker can easily obtain representative datasets released for research purposes on the Internet. Moreover, because the attacker is a participant in the FL system, we consider he may even collude with others to obtain this auxiliary dataset.
Attack Scalability: Our attack is a single-round attack and we have already demonstrated that it can efficiently reconstruct hundreds of samples simultaneously. However, we argue that the attacker can boost his attack scalability performance to launch MediLeak in multiple or continuous FL training rounds. In this way, the attacker can “harvest” the local training samples of a single or even multiple clients continuously to obtain more sensitive data samples. The recovery rate of certain images in the local datasets will also increase, as they may eventually be reconstructed during the continuous multi-round attacks even though they escape the recovery process of a single round.
@@
Defense Mechanisms: One intuitive but effective defense against our attack is to let the clients proactively check the consistency of the model parameters and architectures during the FL training process, rather than blindly believe in the privacy guarantee of the FL systems and give full trust in the FL service providers, which unfortunately, in the current fact for the medical FL systems. However, we argue that our attack can be launched in the initial training rounds or even just the first training round to preserve attack stealthiness. This is because in the initial training rounds, everything is randomly initialized and it is difficult for the defender to distinguish the malicious behaviors from randomly initialized benign patterns.
Other Medical Records: In this work, we focus on the reconstruction problem of the medical images within the secure FL systems under the protection of advanced cryptography mechanisms, demonstrating a serious privacy threat to the current systems. But we notice that the other formats of medical records such as tabular and human language data also contain a huge amount of sensitive and private information. In practice, these records usually involve more detailed personal information such as the patient’s name, address, phone number, and health conditions. We consider the privacy issue of these data records, particularly whether these records can be reconstructed and exposed to a malicious party when they are used in a federated learning system as a very important security and privacy problem. We consider our method to provide a fundamental idea and attack flow about how to recover sensitive client-owned records from the aggregated model updates. However, we acknowledge that such a complex reconstruction problem involving multi-modality data records is non-trivial and deserves a lot of efforts. We leave this as the future direction of our work.
VII Conclusion
In this paper, we present MediLeak–a powerful MIA that breaks the privacy-preserving feature of the current secure federated learning systems protected by the state-of-the-art SA protocols by accurately and efficiently recovering local training samples owned by the clients, resulting in a complete exposure of the patient’s private information. Our attack requires the attacker to first modify the global model architecture and initialize it to form different adversarial modules before publishing it to the clients. Then according to normal FL procedure, the clients train the received global model with their local datasets to provide model updates for the server. Upon recipient, we devise an efficient analytical method to reverse the collected aggregated model updates back to the local training samples without relying upon any costly optimization methods. We implement our attack on two practical medical image datasets and evaluate the attack performance under different settings. The experiment results demonstrate that our attack achieves excellent reconstruction performance, and outperforms the existing optimization-based MIAs. Our attack represents a practical vulnerability of the current medical FL systems, making the community rethink the privacy guarantee of the current systems, and calling for effective defense against such kinds of advanced MIAs.

### v2

@@ 5.2. Image Reconstruction Results
In Tab. 2, we demonstrate the performance of MedLeak over different recovery batch sizes (i.e. the number of samples recovered simultaneously held by the target victim) ranging from 100 to 500 images. We can observe that both the recovery rate (i.e. the ratio of successfully recovered images) and the quantitative scores (i.e. the PSNR and SSIM scores) decrease when the recovered batch size increases. This is expected because the larger the recovery batch size, the more difficult the recovery task to conduct. But in general, MedLeak achieves high recovery rates ($>0.8$) and quantitative scores (PSNR$>80$ and SSIM$>0.9$) for all three datasets under all recovery batch sizes. Particularly, the SSIM scores (ranging from 0 to 1) remain to be 0.99 for all settings, because of the recovery excellency. This can be further verified by the recovered samples we visualized in Fig. 4, in which we plot the original images on the left and the recovered ones on the right. We find that the recovered images are of high quality and cannot be visually distinguished from the original ones, even for some detailed small marks and notations. In Tab. 2 we also demonstrate the attack time (in seconds). We find that the attack time is monotonically increasing with respect to the recovery batch size. For the largest batch size (i.e. 500 images) over the complex COVIDx CXR-4 dataset, it only takes the attacker less than 10 seconds to fulfill the recovery task, indicating that MedLeak is very effective.
5.3. Benchmark Comparison
@@
We compared MedLeak’s attack performance with three optimization-based model inversion attacks (MIAs) including the DLG/iDLG (Zhu et al., 2019; Zhao et al., 2020), InvertGradient (Geiping et al., 2020), and GradInversion (Yin et al., 2021) attacks (denoted as DLG, Invert, and GradInvert respectively) over the MedMNIST dataset for one small batch of input images. We only focused on the image recovery task because all the existing attacks cannot be adapted to the text recovery task. We compared the PSNR scores, SSIM scores, and the attack time between the three attacks and our work. The results are demonstrated in Fig. 5. We can find that our attack achieves much better PSNR scores and SSIM scores than the existing MIAs, indicating that our attack can reconstruct samples with better quality. At the same time, our attack consumes significantly less time than the current MIAs. Particularly, the existing attacks consume a few hundred seconds to reconstruct one batch of samples, while our attack only requires less than one second, which reduces the current cost by two orders of magnitude. The reason why our attack is much more efficient is that our attack only involves closed-form mathematical calculations while the other three attacks require costly iterative-based optimization methods. However, there is no free lunch and we clarify that the three benchmark works adopt an honest-but-curious attack model, which does not allow the attacker to modify the model parameters and architecture as we did.
Table 4. The text reconstruction performance of MedLeak over a different number of text samples and experiment settings. The rate (sample recovery rate) is on a scale of 1.00. The “Embed Dim” refers to the embedding dimension.
Text Num Max Length Embed Dim Rate WER Time (in sec)
20 200 words 64 0.9375 0.0004 0.2149
300 words 64 0.9669 0.0009 0.3057
40 200 words 64 0.9212 0.0004 0.416
300 words 64 0.9153 0.0018 0.6023
60 200 words 64 0.8729 0.0005 0.6341
300 words 64 0.9083 0.002 0.9192
80 200 words 64 0.8228 0.0023 0.8331
300 words 64 0.8540 0.0051 1.2308
100 200 words 64 0.755 0.0047 1.058
300 words 64 0.7585 0.0052 1.514
(a) Recovery bins factor.
(b) Client number factor.
(c) Local training epochs factor.
(d) Non-iid factor.
Figure 6. The recover performance of MedLeak over different practical attack factors.
5.4. Vision Downstream Tasks
@@
To further evaluate the performance of our attack on clinically relevant downstream tasks, we performed a binary disease classification (the detection of COVID-19) task on both the recovered samples and the actual samples. We used the state-of-the-art vision transformer model (ViT-S) (embedding size=368, number of heads=6, 22M parameters) pre-trained by self-supervised learning (SSL) technique on 30k COVIDx CXR-3 samples and fine-tuned on the RSNA-RICORD part of the dataset to perform the classification task (Anwar et al., 2023) and evaluated it on the COVIDx CXR-4 dataset. We demonstrate the performance in Tab. 3. We use widely used machine learning metrics to evaluate the classification performance and we find that the recovered images achieve nearly the same performance as the original ones. This shows that our reconstruction process is highly successful in keeping all semantic meaning within the images and the reconstructed images can be used to perform any potential clinical analysis, which further indicates the severity of the privacy threat imposed by our attack. We consider it a very practical attack scenario for a curious party, which either can be the medical federated learning’s participants or a third-party service provider (who provides the necessary platform, computation resource, and other FL infrastructures) to launch our attack to first reconstruct the sensitive medical images and then feed them to certain downstream analysis tasks to obtain further information about of the victims.
5.5. Text Reconstruction Results
In Tab. 4, we demonstrate MedLeak’s performance on recovering medical text data under different batch sizes (ranging from 20 to 100) and data settings. More specifically, “text num” refers to the number of text samples recovered simultaneously in one batch, and “max length” refers to the maximum length of the text contents in the number of words. In our experiment, we fixed the length of each text sample for processing convenience. We truncated the samples when their lengths were longer than the maximum length and padded them when they were shorter. We also fixed the embedding dimension as the commonly used value 64. From the result, we can find that in general, MedLeak achieves decent text recovery performance under different settings to obtain high recovery rates ($>0.75$), low WER scores ($<0.006$), and short execution time within a few seconds. We observe that when the recovery batch size increases, the recovery rate decreases accordingly, meaning that recovering a larger batch of samples is more difficult. By comparing MedLeak’s performance under different sample lengths (i.e. 200 v.s. 300 words), we further observe that a larger sample length triggers longer attack time, which approximately follows a linear relationship with the sample length. However, a larger sample length does not trigger any recovery performance drop. We still observe decent and stable recovery rates and WER scores when the sample length increases. In fact, the recovery rate even increases when the sample length is larger. This may be because longer samples maintain more semantic information and can be better separated and recovered.
5.6. Performance Affecting Factors
In this section, we investigate MedLeak’s performance under different FL settings. We implemented the attack on the COVID CXR-4 dataset and evaluated the recovery rates and PSNR scores considering the following 4 factors: recovery bin number, client number, local training epoch, and non-iid data settings.
Recovery Bin Number: As we have discussed in Section 4. The recovery bin size $k$ (i.e. neuron number of the first linear layer) can significantly affect the performance of MedLeak. In the experiment, we fixed the reconstruction batch size to 100 and changed the recovery bin $k$ size from 50 to 250. We demonstrate the results in Fig. 6(a). We find that both the recovery rates and PSNR scores are monotonically increasing with $k$ and obtain relatively high quantitative scores when $k\geq m$. This is consistent with our theoretical analysis as more bins (larger $k$) will decrease the recovery conflict probability and increase the recovery success rates. In practice, the attacker can adjust the bin size according to local sample size to obtain decent attack performance.
Client Number: We increased the client number from 5 to 25 in our experiment. The results are shown in Fig. 6(b). We find that the attack performance is not affected by the FL client number. This indicates that our dis-aggregation attack phase is highly successful and the attacker can always obtain accurate model updates of the single victim client.
Local Training Epoch: We increased the local training epochs of the FL clients from 1 to 9. The results are demonstrated in Fig. 6(c). From the results we only observe very slight performance degradation when the FL clients conduct more local training epochs, demonstrating that our attack can be applied to both FedSGD and FedAVG settings.
Non-iid Data Distribution: We changed the class of samples held by the victim clients in our experiment including having COVID-only, non-COVID-only, and both type of samples. The results are shown in Fig. 6(d). We observe that the attack performance is not affected by the type of samples held by the victim client, indicating that MedLeak are applicable to both iid and non-iid settings.
6. Discussion and Future Work
@@
Auxiliary Dataset: Using an auxiliary dataset is a commonly used prerequisite for state-of-the-art MIAs (Pasquini et al., 2022; Fowl et al., 2021; Zhao et al., 2023; Shi et al., 2025). Our attack takes the same assumption and its performance is also affected by the number and quality of samples in the auxiliary data $D_{aux}$. In the ideal case, the auxiliary dataset shall have the same data distribution as the target victim’s local dataset, or the auxiliary dataset is more representative. It may be a challenge for the general vision tasks to find such a representative auxiliary dataset. However, for the medical data, this does not pose a critical barrier, since the radiology data (such as the CT scans) of humans are acquired in similar data format with common anatomical features. The attacker can obtain representative datasets released for research purposes in public domains. Moreover, because the attacker is a participant in the FL system, we consider they may even collude with others to obtain this auxiliary dataset.
@@
Defense Mechanisms: An intuitive yet effective defense against our attack is for clients to proactively verify the consistency of model parameters and architectures during the FL training process, rather than blindly trust the privacy guarantees of FL systems and place full confidence in FL service providers—a trust that is prevalent in the current medical FL systems. However, we argue that our attack can be initiated during the initial training rounds, or even in the very first round, to maintain its stealthiness. This is because, in the early stages of training, when everything is randomly initialized, it becomes challenging for defenders to distinguish between malicious activities and benign patterns that arise from random initialization.
Another potential defense mechanism comes from the data synthesis perspective, leveraging the inherent bottleneck of MedLeak. Our comprehensive analysis reveals that the recovery bin size, denoted as $k$, plays a crucial role in affecting the attack performance. Specifically, when the sample size $m$ is much larger than $k$, MedLeak’s performance drops significantly. Based on this finding, the clients can generate large mask sets containing many images but not exposing anything related to the original private local samples to crowd the recovery bins. When the size of the mask set is large enough, there will be a lot of collisions within the bins, and the attacker can hardly recover anything. This defense strategy can also be employed for other attacks that face similar performance bottlenecks such as (Fowl et al., 2021; Zhao et al., 2023). However, the primary design challenge for such defense is to avoid causing any performance degradation when these mask sets are involved in the FL training process.
More than Linear Leakage: In this work, we leverage the fundamental “linear leakage” primitive as a powerful mathematical tool to help us accomplish our attack. However, we acknowledge that other types of model components such as the vision transformer (Lu et al., 2022), and convolutional layers (Zhao et al., 2023) may also be exploited to reverse the model updates and leak private training data. These model components are integral to popular machine-learning models and can be exploited to launch the attack without the need to insert any additional malicious modules. This would make the attack more stealthy and adaptable for different victim models. However, designing such an analytical gradient reverse method is non-trivial, and we intend to explore such designs in future work.
7. Conclusion
In this paper, we present MedLeak–a novel MIA that targets current FL systems designed for healthcare applications using sensitive patient data. MedLeak can accurately and efficiently recover local training samples at a clinical site, resulting in unwanted leakage of private patient information. To achieve this, MedLeak requires the parameter server (i.e., the attacker) to actively craft additional adversarial attack modules before the global models. These adversarial models are designed with the mathematical guarantee to effectively break the secure aggregation protocol and efficiently recover hundreds of samples in a batch without relying on costly optimization methods when they are sent to the clients. We customize MedLeak to recover both image and textual data records, as clinical data usually comprises both types of samples, extending its applicability to wider healthcare-related FL systems. We implement MedLeak on multiple medical images and text datasets and our results highlight MedLeak’s excellent attack performance under various real-world settings. Our attack exposes a practical vulnerability of the current medical FL systems, prompting the community to reconsider the privacy guarantees of these systems and to develop effective defenses against such advanced MIAs.
Acknowledgments
This work was supported in part by the Office of Naval Research under grants N00014-24-1-2730 and N00014-19-1-2621, the National Science Foundation under grants 2312447, 2247560, 2154929, 2332675, and 2235232, and Children’s National Hospital, the Sanghani Center for AI and Data Analytics, and the Fralin Biomedical Research Institute at Virginia Tech.